CH 16-16
Beckheart is seeking financing for its inventory. Safe-proof Warehouses offers space in their facility for Beckheart’s inventory.
They offer loans with a 15 percent APR equal to 60 percent of the inventory. Monthly fees for the usage of the warehouse are $500 plus 0.5 percent of the inventory’s value. If Beckheart has saleable inventory of $2 million,
A. how much money can the firm borrow?
B. what is the interest cost of the loan in dollars over a year?
C. what is the total amount of fees to be paid in a year?
D. what is the effective annual rate of using Safe-proof to finance Beckheart’s inventory?
1
Expert's answer
2012-10-10T08:56:42-0400
1. Firm borrow = $2 mln*0.6 = $1.2 mln
2. Interest cost = 0.15*1.2 mln =
$0.18 mln
3. Fees = $500+ 0.005*$2 mln = $10500
4. EAR = $0.18 mln/ $1.2 =
0.15 (15%)